# Forcing Planets to Evolve: The Relationship Between Uranus and Neptune at Late Stages of Dynamical Evolution

### v1

@@ 3 One and two planet examples
Figure 3: Fifty million year evolution of Neptune’s inclination for three different simulations initialized as described by column one in Table 2 except Neptune’s inclination is $20^{\circ}$. We artificially evolve Neptune’s inclination as a sinusoid, varying $\Delta i$ and $\tau_{i}$ relative to the secular amplitude, $\mathcal{A}_{\rm sec}$ and oscillation timescale, $\tau_{\rm sec}$ with three examples: (1) $\Delta i>\mathcal{A}_{\rm sec}$ and $\tau_{i}>\tau_{\rm sec}$ (top panel), (2) $\Delta i>\mathcal{A}_{\rm sec}$ and $\tau_{i}<\tau_{\rm sec}$ (middle panel), and (3) $\Delta i<\mathcal{A}_{\rm sec}$ and $\tau_{i}>\tau_{\rm sec}$ (bottom panel).
The artificial evolution of $\omega$ and $\Omega$ in a multi-planet case requires a more careful consideration than that of the other elements. One cannot choose completely arbitrary functions for their evolution because the interactions between the planets dominate many attempts to control them. If planets are not strongly coupled, then the phase angles will be easier to force, but this is typically not the case in a multi-planet system. For example, for several test simulations with Jupiter and Neptune, we added and subtracted a small amount of precession to Jupiter by imposing a linear evolution on top of the $\Omega_{J}$. This is helpful if one desires to slow down or speed up the precession of a planet. In these simulations, Neptune’s $\Omega_{N}$ also increased or decreased, respectively, such that $\Omega_{J}-\Omega_{N}$ remained constant. In another test (not shown), we increased $\Omega_{J}$ and decreased $\Omega_{N}$ so that $\Omega_{J}-\Omega_{N}$ changed over time. For this test, the amplitude and frequency of the eccentricity and inclination varied over time in combination with the change of $\Omega_{J}-\Omega_{N}$, thus changing the secular structure. One example of increasing Jupiter’s $\omega_{J}$ is shown in Figure 4. Ficticious evolution of the phase angles is worth considering for scenarios where the user is interested in altering the interactions between the planets, for example in order to track how varying precession affects thousands of particles in the disk. We discuss these implications further in Section 5.
@@
Figure 4: Fifty million year evolution of Jupiter’s $\omega$ (orange) under the gravitational influence of a solar-mass star and Neptune-mass planet. This simulation is initialized with the parameters described in column one of Table 2. In another simulation, we add a linear perturbation over time to Jupiter’s $\omega$ (maroon line), making Jupiter precess faster (blue line).
@@ 4 Applications for Artificial Planetary Evolution
Here we show appropriate applications for which to use this code. In planetary instability models (e.g. Tsiganis et al., 2005; Levison et al., 2008, 2011; Nesvorný, 2018; de Sousa et al., 2020), the rearrangement and migration of the outer Solar System planets is typically ignited by Jupiter and Saturn crossing or exiting from a mutual mean motion resonance, which causes their orbital eccentricities to be excited. The resulting planetary instability excites the orbits of the other giant planets, and they undergo a chaotic era of strong planet-planet interactions for several millions to tens of millions of years. In simulations with successful outcomes that qualitatively match the real solar system, Uranus and Neptune, characterized by their eccentric and inclined orbits, are typically damped to more circular and coplanar orbits by dynamical friction (i.e. exchange of angular momentum and energy between the planets and nearby planetesimals). At this time, Jupiter and Saturn were most likely damped to some extent as well but not as much as the outer ice giants. The details of when an instability occurred and how long or how far Neptune migrated remain uncertain. The unprecedented outer solar system de-biased survey made by the Vera C. Rubin Observatory Legacy Survey of Space and Time will help constrain the giant planets evolution even further by allowing comparisons betweeen simulations and observations. In the context of instability models, we can split the possible story of the outer solar evolution into two phases: a chaotic phase and a damping (and migration, though we focus on eccentricity damping) phase.
The behavior of the planets in the chaotic era is unpredictable. Simulations with minuscule changes in the initial conditions of a planetary system produce vastly different planetary behaviors and outcomes. The original planetary instability model (the Nice model—Tsiganis et al. (2005)) included a fifth giant planet that was expelled from the system, and the two farthest planets switched locations in a fraction of simulations. The stochasticity of planet behaviors for various initial conditions makes statistical validation of the initial conditions and outcomes challenging. In addition, massive planetesimals are necessary in the simulation for two important aspects of this model: (1) instigating the initial mean motion resonance crossing between Jupiter and Saturn that leads to an instability and (2) damping Uranus’ and Neptune’s eccentricity which facilitates a stable configuration after tens of millions of years. Including massive planetesimals in such a simulation, makes this a more computationally expensive problem. The interactions between the planets and planetesimals causes the simulations to run for a longer amount of time and adds an extra layer of chaos. To mitigate this complication, artificial forces—as discussed in this paper—can be used to mimic the presence of a massive planetesimal disk and artificially evolve planets’ orbits.
In this section we show two example applications of artificial planet evolution: (1) reproducing phases of migration and eccentricity damping after a planetary instability and (2) investigating the secular coupling between Uranus and Neptune by artificially damping Neptune’s eccentricity and seeing the effect on Uranus’ eccentricity. For both applications, our initial conditions are set by the planet parameters after a chaotic evolution/before a migration from the simulation described in Figure 1 of Tsiganis et al. (2005).
@@ 4.1 Reproducing Epochs of Migration and Eccentricity Damping
We extrapolate the eccentricity evolution of the giant planets from Figure 1 in Tsiganis et al. (2005) using the semi-major axis, apocenter, and pericenter curves. [^1] We use the online tool Webplot Digitizer to extract the data from the figure (https://apps.automeris.io/wpd4/). We calculate "early" and "late" eccentricity time-evolution profiles for Uranus and Neptune using the values in the three curves since $a_{\rm peri}=a(1-e)$ and $a_{\rm apo}=a(1+e)$. (See Figure 5). We fit an exponential function $\Delta e+e_{0}+\Delta e\exp{t/\tau}$ to the data to find a best-fit damping timescale, $\tau$. The presence of a massive planetesimal disk in Tsiganis et al. (2005) simulations is responsible for damping the eccentricities of Uranus and Neptune. After millions of years, the volume density of the disk can drop by 2-3 orders of magnitude, so the damping rate decreases as well. Our goal here is to calculate damping timescales due to analytical dynamical friction theory and compare with the best fit damping timescales extracted from the simulations to confirm dynamical friction is the dominant force at play. With this example, we will demonstrate how one can artificially evolve planets according to a physically motivated process.
To calculate an eccentricity damping timescale due to dynamical friction we use the parameters of the disk stated in Tsiganis et al. (2005) and in the caption of their Figure 1. The information that is not explicitly stated is assumed to the best of our ability and does not impact the timescale substantially. The initial orbits of Jupiter, Saturn, Uranus, and Neptune are nearly circular and coplanar ($e\sim i\sim 0.001$) with initial semimajor axes 5.45, 8.65, 17, and $\sim$11, respectively. In that simulation, Uranus and Neptune switch orbital positions, a phenomenon that occurred $\sim 50$% of the time. Their simulated planetesimal disk is 35 $M_{\rm E}$, composed of 3500 equal-mass particles up to 30 au, and has a surface density that falls linearly with heliocentric distance. We assume an inner edge of 17 au, since it is not stated in the paper. All particles are initialized with ($e\approx\sin i\approx 0.05$). These values are essential in calculating the timescale at which Neptune would damp due to dynamical friction, since the amount of momentum exchange between the planets and the planetesimals in the disk depends on the volume density of the medium.
Encounters between the planet and sea of planetesimals changes the relative velocity between the two bodies, which is the difference between the orbital velocity of the planet and mean disk flow. The relative velocity can be dominated by the planet’s or planetesimal’s random velocity, which we define as $v^{2}=v_{l}^{2}+v_{P}^{2}$, where $v_{l}$ is the planetesimal’s random velocity and $v_{P}$ is the planet’s random velocity. The order of magnitude momentum exchange of the encounter per time is given by $M_{P}\frac{dv}{dt}=-(mv)n\sigma v_{=}-\rho\sigma v^{2}$, where $M_{P}$ is the mass of the planet, $m$ is the mass of an individual planetesimal, $\rho$ is the volume density of the disk, $\sigma$ is the cross section area for the encounter. Assuming the strong scattering cross section determined by the impact parameter $b\approx GM_{P}/v^{2}$, the equation simplifies to $M_{P}\frac{dv}{dt}=-\rho\Big{(}\frac{GM_{P}}{v^{2}}\Big{)}^{2}v^{2}$. This form matches that found in Goldreich et al. (2004) and Ford & Chiang (2007), which we follow for a portion of this calculation. We adopt constants that were dropped in the order of magnitude expression moving forward. The dynamical friction force is thus given by
@@
Figure 5: Eccentricity for Uranus (top panel) and Neptune (bottom panel) calculated by using the semimajor axis, pericenter, and apocenter curves from Figure 1 of Tsiganis et al. (2005). The best-fit exponential damping timescale at the early stage (before 20 Myr) is 1.5 Myr for both Uranus and Neptune. At late times (after 30 Myr), the best-fit exponential timescales are 20 Myr and 6 Myr for Uranus and Neptune, respectively (see Section 4.1).
$\frac{dv}{dt}=-\frac{2\pi G^{2}M_{P}\rho}{v^{2}}\ln(1+\Lambda^{2})$ (8)
where the Coloumb parameter, $\Lambda$ is
$\Lambda=b_{\rm max}\big{(}v^{2}+2\langle\sin(i_{l})^{2}\rangle v_{\rm circ}^{2}\big{)}\Big{/}(GM_{P})$ (9)
and
$b_{\rm max}=R_{H}+r\big{(}\langle\sin(i_{l})^{2}\rangle+\sin(i_{P})^{2}\big{)}^{1/2}$ (10)
The subscripts $l$ and $P$ represent the planetesimal and planet, respectively. The calculation in Ford & Chiang (2007) assumes the planet interacts with the disk twice since they assume $i_{P}\gg i_{l}$, whereas we are assuming that the planet and planetesimal disk have similar inclinations (i.e. the planet is embedded in the disk and feels the perturbation on behalf of the disk for the duration of an orbital period). The change in planet’s velocity due to the interactions with the planetsimals is $\Delta v_{P}$ and the interaction time, $\Delta t$, is the orbital period, $2\pi/\Omega$. The kicks thus produce $\Delta v_{P}\approx\frac{dv}{dt}\Delta t$, so that
$\frac{dv_{P}}{dt}=\frac{\Delta v_{P}}{t_{\rm orb}}=-\frac{2\pi G^{2}M_{P}}{v^{2}}\ln(1+\Lambda^{2})\rho$ (11)
$\frac{de_{P}}{dt}a\Omega\hat{\boldsymbol{p}}=-\frac{2\pi G^{2}M_{P}}{(\sqrt{2}e_{P}a\Omega)v^{2}}\ln(1+\Lambda^{2})\rho(e_{P}a\Omega\hat{\boldsymbol{p}}+i_{P}a\Omega\hat{\boldsymbol{z}})$ (12)

### v2

@@ 3 One and two planet examples
Figure 3: Fifty million year evolution of Neptune’s inclination for three different simulations (solid lines) initialized as described by in Table 1. We artificially evolve Neptune’s inclination (dashed lines) as a sinusoid, varying $\Delta i$ and $\tau_{i}$ relative to the secular amplitude, $\mathcal{A}_{\rm sec}$ and oscillation timescale, $\tau_{\rm sec}$ with three examples: (1) $\Delta i>\mathcal{A}_{\rm sec}$ and $\tau_{i}>\tau_{\rm sec}$ (top panel), (2) $\Delta i>\mathcal{A}_{\rm sec}$ and $\tau_{i}<\tau_{\rm sec}$ (middle panel), and (3) $\Delta i<\mathcal{A}_{\rm sec}$ and $\tau_{i}>\tau_{\rm sec}$ (bottom panel).
@@
Figure 4: Two, fifty million year evolutions of Jupiter’s $\omega$ under the gravitational influence of a solar-mass star and Neptune-mass planet. Both simulations are initialized with the parameters described in Table 1. The first simulation does not include any additional velocities or accelerations (orange). In the second simulation, we add a linear perturbation (maroon) over time to Jupiter’s $\omega$, making Jupiter precess faster (blue).
For the two-planet example shown in Figure 2, we simulate a Jupiter-mass planet and Neptune-mass planet for one hundred and fifty million years. The planets are initialized with values and imposed evolution shown in Table 1. Only the semimajor axis, eccentricity, and inclination are artificially evolved. As illustrated by the oscillation in the eccentricity and inclination, including a second planet in the system introduces gravitational forces between the planets which were not present in the one-planet simulation shown in Figure 1. In this three-body system, the planets secularly perturb each other, causing oscillations in the eccentricity and inclination with periods shorter than the damping timescale we prescribe. In the one-planet example, the orbital elements follow the prescribed evolution perfectly, but that’s not the case here. Rather than having full control of the final value of a given element, the scale of change, $\Delta g$ (i.e. final value minus initial value), in Equation (7c) sets an approximate final outcome in simulations with more than one planet. Jupiter’s secular effect on Neptune is greater than Neptune’s effect such that Neptune often over- or under-shoots the prescribed final eccentricity by an amount comparable to the amplitude of the secular oscillation. We find that when prescribing an exponential growth or decay for eccentricity or inclination, the simulated average $e$ or $i$ may asymptote toward a value that is greater or smaller than the specified final value, as in Figure 2. In analyzing hundreds of simulations, we did not find an obvious trend for how simulation conditions affect the final simulation $e$ or $i$ value relative to the prescribed final value.
While the average simulation $e$ or $i$ value may asymptote below or above the value we prescribe, this is not a major issue because the simulation is still evolving accurately under the conditions provided. This could, however, be an issue if the user’s primary goal is to reach a specific average value. Damping to an exact value will not happen each time in the 2+ planet simulations because the gravitational perturbations between the planets may dominate under certain conditions. This interdependence is important because we are interested in modeling how planets behave under physical (absent from the simulation) processes in addition to planet interactions. We note that because we are using additional accelerations and velocities to model the behavior of an absent physical process, we are not conserving the angular momentum or energy of the planet. The energy and angular momentum gained or lost by the planet would be lost or gained by the population of planetesimals (or other bodies) that the fictitious forces and velocities are mimicking.
The simulation setup in Figure 2 is such that Neptune’s secular inclination amplitude is large since Jupiter started at a high inclination. The secular eccentricity perturbations for both Jupiter and Neptune are small compared to the eccentricity user-defined perturbations (i.e. the total user input change is several times larger than the secular amplitude), whereas, the exponential change for the inclination is on the same order as the amplitude of the secular oscillations. Understanding how user-defined perturbations compare to the gravitational perturbations between planets is essential for running effective models and knowing the limit of control one has on the evolution of planets.
We highlight the limitation for controlling planets’ evolution further in Figure 3 with three different simulations. Each simulation includes the star and two planets with parameters defined in Table 1. We apply a sinusoidal evolution to both planets’ inclination to illustrate how the evolution differs for different values of $\Delta i$ and $\tau_{i}$ relative to the secular amplitude, $\mathcal{A}_{\rm sec}$ and secular oscillation timescale (1/frequency), $\tau_{\rm sec}$. Figure 3 shows the Neptune-mass planet’s inclination evolution over fifty million years. When $\tau_{i}>\tau_{\rm sec}$, the long-term evolution is dominated by the user-defined forces. When $\Delta i>\mathcal{A}_{\rm sec}$, the prescribed sinusoidal evolution is noticeable, as demonstrated in the top panel of Figure 3. The next two simulations show the case where $\tau_{i}<\tau_{\rm sec}$ and $\Delta i<\mathcal{A}_{\rm sec}$, respectively. When $\Delta i<\mathcal{A}_{\rm sec}$, the secular oscillation frequency is equivalent to the case when no perturbations are present. Note, when the amplitude of the induced inclination oscillation is higher (top panel; Figure 3), the secular oscillation frequency is lower compared to the case with a small amplitude imposed oscillation (bottom panel). The secular frequency is altered because the imposed evolution is sinusoidal; we are effectively adding a frequency that interferes with the natural secular oscillation frequency. If we instead impose an exponential function, which is more typical for damping and exciting eccentricities and inclinations, the secular frequency does not substantially change (not shown).
The artificial evolution of $\omega$ and $\Omega$ in a multi-planet case requires a more careful consideration than that of the other elements. One cannot choose completely arbitrary functions for their evolution because the interactions between the planets dominate many attempts to control them. If planets are not strongly coupled, then $\omega$ and $\Omega$ will be easier to force, but this is typically not the case in a multi-planet system. For example, for several test simulations with Jupiter and Neptune, we added and subtracted a small amount of precession to Jupiter by imposing a linear evolution on top of the $\omega_{J}$. This is helpful if one desires to slow down or speed up the precession of a planet. In these simulations, Neptune’s $\omega_{N}$ remained the same, and $\varpi_{J}-\varpi_{N}$ remained constant, where $\varpi=\Omega+\omega$. In another test, we increased $\omega_{J}$ and decreased $\omega_{N}$ so that $\varpi_{J}-\varpi_{N}$ changed over time. For this test, the amplitude and frequency of the eccentricity and inclination of each planet varied over time, thus changing the secular structure. One example of increasing Jupiter’s $\omega_{J}$ is shown in Figure 4. Fictitious evolution of $\omega$ and $\Omega$ is worth considering for scenarios where the user is interested in altering the interactions between the planets, for example in order to track how varying precession affects thousands of particles in the disk. We discuss these implications further in Section 5.
@@ 4 Applications for Artificial Planetary Evolution
In this section we show two example applications of artificial planet evolution: (1) reproducing phases of migration and eccentricity damping for Uranus and Neptune after a planetary instability (Section 4.1) and (2) investigating the secular coupling between Uranus and Neptune by artificially damping Neptune’s eccentricity and seeing the effect on Uranus’ eccentricity (Section 4.2). For both applications, our initial conditions are set by the planet parameters after a chaotic evolution and before a period of smoother migration from the simulation described in Figure 1 of Tsiganis et al. (2005).
@@ 4.1 Reproducing Epochs of Migration and Eccentricity Damping
First, we describe the two phases of semi-major axis and eccentricity evolution from Tsiganis et al. (2005) when fit by an exponential function (Section 4.1.1). Next, we calculate the eccentricity damping timescale due to analytical dynamical friction theory and compare it with the simulation, best-fit timescale to confirm that dynamical friction is the dominant force at play (Section 4.1.2). Lastly, we present two simulations where we artificially evolve the planets’ semimajor axis and eccentricity and show that we reproduce migration and eccentricity damping phases from Tsiganis et al. (2005) (Section 4.1.3). With this example, we will demonstrate how one can artificially evolve planets according to a physically motivated process.
@@
Figure 5: Eccentricity for Uranus (top panel) and Neptune (bottom panel) calculated by using the semimajor axis, pericenter, and apocenter curves from Figure 1 of Tsiganis et al. (2005). The best-fit exponential damping timescale at the early stage (before 20 Myr) is 1.5 Myr for both Uranus and Neptune. At late times (after 30 Myr), the best-fit exponential timescales are 20 Myr and 6 Myr for Uranus and Neptune, respectively (see Section 4.1.1 for the best fit final and initial eccentricities).
4.1.1 Finding the Best-fit Exponential Migration and Eccentricity Damping Timescales
We extrapolate the semi-major axis, apocenter, and pericenter evolution of the giant planets from Figure 1 in Tsiganis et al. (2005) for the time range between 8.6 Myr and 80 Myr. [^1] We use the online tool Webplot Digitizer to extract the data from the figure (https://apps.automeris.io/wpd4/). We calculate “early” (8.6-20 Myr) and “late” (30-80 Myr) eccentricity time-evolution profiles for Uranus and Neptune using the values in the three curves since $a_{\rm peri}=a(1-e)$ and $a_{\rm apo}=a(1+e)$ (see Figure 5). We use scipy.optimize.curve_fit to fit an exponential function (Equation 7c) to the data to find best-fit parameters for semi-major axis and eccentricity. Because the fit is done with data points that were extracted by hand, which are not evenly spaced in time, this fit is just meant to be an estimate of the overall motion. For Uranus’ early and late phases, we find ($e_{0}=0.06$, $e_{f}=0.018$, $\tau_{e}=1.5$ Myr; $a_{0}=18.6$ au, $a_{f}=19.3$ au, $\tau_{a}=7.2$ Myr) and ($e_{0}=0.054$, $e_{f}=0.03$, $\tau_{e}=20$ Myr; $a_{0}=19.3$, $a_{f}=19.7$, $\tau_{a}=17$ Myr), respectively. For Neptune’s early and late damping phases, we find ($e_{0}=0.06$, $e_{f}=0.02$, $\tau_{e}=1.5$ Myr; $a_{0}=23.7$ au, $a_{f}=25.4$ au, $\tau_{a}=11$ Myr) and ($e_{0}=0.055$, $e_{f}=0.022$, $\tau_{e}=6$ Myr; $a_{0}=25.5$ au, $a_{f}=26.43$ au, $\tau_{a}=9.6$ Myr), respectively.
In this simulation from Tsiganis et al. (2005), massive planetesimals are present, and have a large role in damping the planets’ eccentricities through dynamical friction. The damping timescale and amount of damping due to dynamical friction depend on the mass density of the planetesimal disk (i.e. a larger mass in planetesimals per volume produces more momentum exchange between a planet and its neighboring planetesimals, and the planet will damp faster and further). At early times, the mass density across the planetesimal disk is higher than late times and likely comparable at the locations of Uranus and Neptune, so it is expected that they have similar, short damping timescales in this phase (see below). After continuous strong gravitational scattering events, the mass density decreases; whether this change is uniform across the disk in the simulation from Tsiganis et al. (2005) is not explicitly stated. It is possible that the local mass density near Neptune is larger than that around Uranus and as a result, Neptune’s eccentricity is damped further and faster than Uranus’. We investigate this point by running a few dynamical instability simulations and track how the number density across the disk evolves over time. We found that for the first $\sim$10 Myr after a dynamical instability, the number density remains approximately uniform across the locations of Uranus and Neptune, consistent with the similar damping times evident at early times in Figure 5. At later times, the region near Uranus eventually becomes more cleared, but we cannot verify the timescale of that clearing in the Tsiganis et al. (2005) simulation. We note that it is possible that the secular coupling between Uranus and Neptune has a role in this late damping phase. We discuss this point further in Section 4.2.
4.1.2 Calculating Eccentricity Damping Timescale due to Dynamical Friction
To calculate an eccentricity damping timescale due to dynamical friction we use the parameters of the disk stated in Tsiganis et al. (2005) and in the caption of their Figure 1. The information that is not explicitly stated is inferred to the best of our ability and does not impact the timescale substantially. The initial orbits of Jupiter, Saturn, Uranus, and Neptune are nearly circular and coplanar ($e\sim i\sim 0.001$) with initial semimajor axes 5.45, 8.65, $\approx$17, and $\approx$12, respectively. In that simulation, Uranus and Neptune switch orbital positions, a phenomenon that occurred $\sim 50$% of the time. Their simulated planetesimal disk is 35 $M_{\rm E}$, composed of 3500 equal-mass particles up to 30 au, and has a surface density that falls linearly with heliocentric distance. We assume an inner edge of 17 au, since the outer planet is initially located at this distance. All particles are initialized with ($e\approx\sin i\approx 0.05$). These values are essential in calculating the timescale at which Neptune would damp due to dynamical friction, since the amount of momentum exchange between the planets and the planetesimals in the disk depends on the mass density of the medium.
As a planet travels within a sea of planetesimals, the random velocities of the planet, $v_{P}$, and the planetesimals, $v_{l}$, (i.e. the velocity deviating from the orbital velocity of a circular motion) evolve through dynamical interactions between the bodies. The subscripts $l$ and $P$ represent the planetesimal and planet, respectively. We are in the regime where the planet is much larger than the planetesimals so that $M_{P}v_{p}>m_{l}v_{l}$. As a result, the big body must collide with many small bodies before its velocity changes significantly. It is also the case that $v_{P,\rm esc}>v_{l}>v_{P,\rm H}$, where $v_{P,\rm esc}$ is the escape velocity from the planet’s surface and $v_{P,\rm H}$ is the planet’s Hill velocity. As illustrated in (Goldreich et al., 2004), this means that the primary exchange of momentum between the planet and planetesimals is by collisionless gravitational deflections, enhanced by gravitational focusing. Over time, the exchange of momentum leads to dynamical friction, which damps the planet’s eccentricity (see Goldreich et al. 2004 for a review). The planet will stop damping once equipartition of random kinetic energies is reached (i.e. $M_{P}v_{P}^{2}=m_{l}v_{l}^{2}$, which we don’t expect to occur given that $M_{p}\gg m_{l}$) or once most of the planetesimals have been ejected. Encounters between the planet and planetesimals also change the relative velocity between the two bodies, $v_{\rm rel}\sim\sqrt{v_{l}^{2}+v_{P}^{2}}$, which can be dominated by the planet’s or planetesimal’s random velocity. If the planet’s eccentricity is high enough, $v_{P}>v_{l}$, the dynamical friction formulation is slightly different than the case where planetesimals’ eccentricities are more excited and $v_{l}>v_{P}$. Given that our calculations are meant to model dynamical friction in a dynamical upheaval scenario, we account for both cases.
The order of magnitude momentum exchange per time due to encounters between the planet and planetesimals takes the form $M_{P}\frac{dv}{dt}=-(m_{l}v_{\rm rel})n\sigma v_{\rm rel}$, where $mv$ is the momentum transfer per encounter and $n\sigma v$ is the number of encounters per time. Here, $M_{P}$ is the mass of the big body, $m_{l}$ is the mass of a small body, $n$ is the number density (so that $\rho=m_{l}n$ is the mass density) of the planetesimal disk, and $\sigma$ is the cross sectional area for the encounters. Gravitational focusing is important, so the cross section, $\sigma=\pi b^{2}$, is determined by the impact parameter $b\approx GM_{P}/v_{\rm rel}^{2}$.
[…]
When $v_{P}<v_{l}$, we need to account for momentum exchange in front and behind the planet. This results in $M_{P}\frac{dv_{P}}{dt}\approx n\sigma(v_{l}-v_{P})m_{l}(v_{l}-v_{P})-n\sigma(v_{l}+v_{P})m_{l}(v_{l}+v_{P})\sim-\rho\Big(\frac{GM_{P}}{v_{\rm rel}^{2}}\Big)^{2}v_{l}v_{P}$. This form matches that found in Goldreich et al. (2004). Using $v_{\rm rel}\approx v_{l}$, the dynamical friction force in this regime is thus given by
$\frac{dv_{P}}{dt}=-\frac{2\pi G^{2}M_{P}\rho}{v_{l}^{3}}v_{P}\ln(1+\Lambda^{2})$ (8)
We approximate the maximum and minimum impact parameters as $b_{\rm max}=R_{H}+H$ and $b_{\rm min}=GM_{P}/v_{l}^{2}$, where $R_{H}=a_{P}(M_{P}/3M_{*})^{1/3}$ is the Hill sphere radius, $H=v_{l}/\tilde{\Omega}$ is the scale height of the disk, and $\tilde{\Omega}$ is the orbital angular frequency (Stewart & Ida, 2000).
In equipartition, $\boldsymbol{v}$ has two degrees of freedom parallel to the plane and one degree of freedom perpendicular to the plane, so that $e\approx\sqrt{2}\sin{i}$. We approximate $v_{P}=\sqrt{3/2}e_{P}\tilde{\Omega}_{P}a_{P}$ and $v_{l}=\sqrt{3/2}e_{l}\tilde{\Omega}_{P}a_{P}$, where we have made the simplifying assumption that the semi-major axes and hence orbital angular frequencies of the planetesimals and planet are the same. Plugging in for the planetesimal and planet random velocities, Equation (8) simplifies to